# SAMM: Sharded Automated Market Makers

### v1

@@ Abstract
Automated Market Makers (AMMs) are a cornerstone of decentralized finance (DeFi) blockchain-based platforms. They are smart contracts, enabling the direct exchange of virtual tokens by maintaining liquidity pools. Traders exchange tokens with the contract, paying a fee; liquidity comes from liquidity providers, paid by those fees. But despite growing demand, the performance of AMMs is limited. State-of-the-art blockchain platforms allow for parallel execution of transactions. However, we show that AMMs do not enjoy these gains, since their operations are not commutative so transactions using them must be serialized.
We present SAMM, an AMM comprising multiple independent shards. All shards are smart contracts operating in the same chain, but they allow for parallel execution as each is independent. The challenge is that trading in a standard AMM is cheaper if its liquidity pool is larger. Therefore, we show that simply using multiple smaller AMMs results in traders splitting each trade among all AMMs, which worsens performance. SAMM addresses this issue with a novel design of the trading fees. Traders are incentivized to use only a single smallest shard. We show that all Subgame-Perfect Nash Equilibria (SPNE) fit the desired behavior: Liquidity providers balance the liquidity among all pools, so the system converges to the state where trades are evenly distributed.
Evaluation in the Sui blockchain shows that SAMM’s throughput is over fivefold that of traditional AMMs, approaching the system’s limit. SAMM is a directly deployable open-source smart contract, allowing trading at scale for individuals and DeFi applications.
1 Introduction
@@
Figure 1: Trade transaction latency as a function of demand with $n$ SAMM shards.
@@
In this work, we address the throughput limitation of AMMs by using multiple AMM instances called shards. We model the system (§3) as a set of AMM shards and rational users of two kinds. Traders purchase tokens, they use the available AMMs and pay fees as required, aiming to minimize their expenses. Liquidity providers deposit tokens into AMMs and earn fees based on their contribution.
@@
The shards are AMMs based on the standard Constant Product Market Maker (CPMM) contract (§4). Roughly, the contract maintains the product of the two tokens constant after each trade. Thus, purchasing a larger amount of a token increases its unit cost, an effect called slippage.
We present SAMM (§5), an AMM protocol that uses multiple shards, all of which are AMM smart contracts operating on the same blockchain. Ideally, the shards should be balanced, i.e., have equal liquidity (deposited amounts), and traders should randomly select a shard to complete each trade.
@@
The model gives rise to a game (§6) played among the users. In each step, either a liquidity provider adds liquidity to a subset of the shards, or a trader executes a trade using a subset of the shards. We assume myopic liquidity provider behavior, reducing the analysis to a Stackelberg game where the liquidity provider adds liquidity to maximize her revenue from a subsequent trade. We observe that naively using a set of independent CPMMs results in all trades being split among all CPMMs, increasing system overhead without improving throughput. To overcome this, rather than using a set fee ratio (as in all previous work we are aware of), we set a range of possible fees. Within this range, SAMM uses a trading fee function that encourages traders to use the smallest shard.
Our analysis (§7) shows that, indeed, in all best responses, traders use one of the smallest pools. This, in turn, implies that not filling the smallest pool is not the best response for a liquidity provider. We provide specific strategies for traders and liquidity providers that form a subgame perfect equilibrium. We also show that, once the system reaches the balanced state, it will stay in that state.
@@
To evaluate SAMM (§8), we implemented the protocol and deployed it to a local test network of the Sui blockchain platform [10]. SAMM achieves over a fivefold throughput increase compared to a standard single-contract AMM. Figure 1 shows that with more shards, SAMM achieves higher throughput (X axis) with lower trimmed-mean latency (Y axis). Error bars show additional experiments, X marking failure due to overload. This increase is limited by the serial elements of Sui’s transaction processing, following Amdahl’s Law. Finally, we confirm the theoretical analysis by simulating trades from real data and observe that (1) traders follow the desired behavior and (2) SAMM significantly improves the liquidity providers’ revenue with a minor increase in the traders’ costs due to enhanced throughput that allows for more trades.
@@
In summary, our contributions are: (1) Identification of the performance challenges due to “hot” contracts, (2) generalization of the trading fee function of AMMs, (3) SAMM: sharded AMM contract with a novel trading fee function to incentivize the desired behavior, (4) game-theoretic analysis showing Subgame-Perfect Nash Equilibrium (SPNE), (5) evaluation in Sui, demonstrating a fivefold increase in throughput (up to the blockchain’s limit), and (6) simulation with real trade data confirming the theoretical analysis.
@@
These results hint at an upcoming challenge (§9) in smart contract platform design: minimizing the serial elements of transaction processing. But SAMM can already be employed to scale AMM performance both for direct usage and as part of DeFi smart contracts.
2 Related Work
@@
The introduction of the constant product market maker (CPMM) model by Uniswap v1 [52] set a new standard for AMMs, employing a liquidity pool and an algorithm designed to keep the product of the token balances constant. This approach enabled asset exchanges without relying on traditional order books. Subsequent iterations, Uniswap v2 [4] and v3 [3], further developed the CPMM model by improving the price oracle mechanism and the returns for liquidity providers, respectively. As a result, the CPMM algorithm has become a benchmark, with many AMMs adopting similar trading mechanisms [34, 44, 15].
@@
Academic research has primarily concentrated on theoretical models, utility optimization, and security issues surrounding AMMs. Angeris et al.[5] expanded the understanding of AMMs by delving into constant function market makers (CFMMs), demonstrating their utility as decentralized price oracles and broadening the CPMM model’s application. Following this, research has increasingly focused on trading utility maximization [6], advanced arbitrage techniques [53, 26, 7, 46], improving liquidity providers’ returns [32, 23], ensuring transaction privacy [14], eliminating Miner Extractable Value (MEV) for fair trades [13, 48, 12], and examining the synergy between blockchain-based AMMs and prediction market mechanisms [40]. To the best of our knowledge, previous work did not address AMM throughput scaling.
@@
3 Model
We abstract away the blockchain details for our analysis and model the system as a set of participants (§3.1) interacting directly with AMMs (§3.2), exchanging tokens. The system progresses in discrete steps (§3.3) and there exists an external market used by arbitrageurs (§3.4) to arbitrage in our system. The model uses a generic AMM, which we will later instantiate based on previous work and with SAMM.
3.1 Participants
There are two types of participants, liquidity providers and traders. Each liquidity provider holds some token A and token B. They aim to increase their holdings. Traders are either AB or BA, based on their goals. Each BA trader occasionally wishes to get a certain amount of token A, and vice versa for AB traders. They have sufficiently many tokens of the opposite type to complete their trade, but they aim to minimize the cost of obtaining the desired tokens. Both liquidity providers and traders can send and receive tokens to and from the smart contracts. We ignore the gas fee of smart contract executions for simplicity.
3.2 Automated Market Makers
@@
  • Trade: The trader sends $I^{A}$ token A (or, alternatively, $I^{B}$ token B) to the contract and gets $O^{B}$ token B (resp., $O^{A}$ token A) from the contract.
@@
We follow prior studies [32, 23] and assume that the trading fee is directly paid to the liquidity providers. Although many practical AMMs (e.g., Uniswap v2[4]) reinvest the trading fees into the liquidity pool, allowing liquidity providers to withdraw more tokens than they deposited as utilities, the trading fee’s impact is negligible relative to the deposited amount. The average ratio of the output amount to the deposited amount is minimal (less than $0.036\%$ as we find in Appendix B), and automated market makers (AMMs) typically charge a low trading fee relative to the gross amount (specifically, $0.3\%$ in Uniswap), so the trading fee’s impact is negligible relative to the amount of deposited tokens. Therefore, our model remains applicable even if trading fees were to be reinvested, given their negligible size.
3.3 System State and Progress
In the system, there are $n$ independent AMM liquidity pools $\textit{pool}_{1},\textit{pool}_{2},\cdots\textit{pool}_{n}$. The system progresses with discrete steps $k=0,1,2,\cdots$ and is orchestrated by a scheduler. In each step, the scheduler randomly selects a participant and this participant executes transactions. The probability of choosing a liquidity provider is $P_{lp}\geq 0$, while the probability of choosing a trader is $P_{t}\geq 0$, with $P_{lp}+P_{t}=1$.
@@ 3.3 System State and Progress
The trader is either a BA trader aiming to obtain token A or an AB trader aiming to obtain token B. The probability of drawing an AB trader (resp., BA trader) is $P_{t}^{\textit{AB}}\geq 0$ (resp., $P_{t}^{\textit{BA}}\geq 0$), with $P_{t}^{\textit{AB}}+P_{t}^{\textit{BA}}=P_{t}$. To avoid repetition, we only show the case of BA traders, the expressions for AB traders are symmetric. The system assigns the BA trader an amount $b^{\textit{BA}}$ token A to obtain following a random distribution $D^{\textit{BA}}$.
3.4 External Market and Arbitrageurs
Following previous work [23, 32], we assume there is an external market providing the price of token A and token B, $p^{A}$ and $p^{B}$, respectively. These prices do not change due to trades; they serve as objective prices for token A and token B.
When someone can get a lower price of tokens in the AMM than the external market, then he can buy tokens from the AMM and sell them in the external market to make profits or vice versa if the price is higher in the AMM; this is arbitrage. Previous work [32, 23, 12, 13] assume active and rational arbitrageurs who can use the external market and always make arbitrages to maximize their utility. We follow the assumption [32] that there are immediate arbitrages without trading fees when the token price in the AMM is different from the external market.
3.5 Our Goal
@@
4 Preliminaries: CPMM
As a baseline for SAMM, we consider the Constant Product Market Maker (CPMM, e.g. [3, 34, 44, 15]), the prominent AMM in real-world applications. It uses a share-based solution to manage liquidity addition and removal operations (§4.1) and keeps the product of the deposited amount of two tokens constant in trade operations (§4.2). Liquidity providers earn revenue from the trading fees of traders (§4.3).
4.1 Liquidity Addition and Removal
Most CPMMs (e.g., [52, 4, 34, 37]) use a fungible share token to manage liquidity addition and removal operations. These tokens represent a liquidity provider’s share in the pool.
When liquidity providers add tokens to an AMM, they receive share tokens which signify their portion of the pool. To recall, $R^{A}$ and $R^{B}$ denote the amounts of token A and token B already deposited in the pool. Similarly, $I^{A}$ and $I^{B}$ represent the quantities of token A and token B that the liquidity provider contributes through a liquidity addition operation. Let $R^{S}$ represent the total amount of all share tokens distributed before the operation. The amount of share tokens acquired by the liquidity provider in this operation, $O^{S}$, is given by:
@@ 4.1 Liquidity Addition and Removal
This term $\min\left\{\frac{I^{A}}{R^{A}},\frac{I^{B}}{R^{B}}\right\}$ signifies the ratio of the input token to the deposited token. The $\min$ function serves to ensure that the ownership accurately reflects the liquidity provider’s contribution relative to the scarcer asset. It prevents situations where a liquidity provider inputs a large amount of a certain token to unfairly obtain a larger share of tokens in the pool.
Liquidity providers have the option to withdraw tokens from the pool with the liquidity removal operation, which takes input share tokens and outputs token A and token B. Let $I^{S}$ represent the amount of input share tokens, and $R^{S}$ denote the total amount of all share tokens referred to the pool before the execution. The amounts of token A and token B withdrawn are $O^{A}$ and $O^{B}$:
@@ 4.2 CPMM Trades
$\Phi^{\textit{net}}_{\textit{CPMM}}(R^{A},R^{B},I^{A},O^{A},I^{B},O^{B})\mathrel{\mathop{:}}=\\ (R^{A}+I^{A}-O^{A})\times(R^{B}+I^{B}-O^{B})\,\,.$
@@
$\Phi^{\textit{net}}_{\textit{CPMM}}(R^{A},R^{B},I^{A},O^{A},I^{B},O^{B})=\Phi^{\textit{net}}_{\textit{CPMM}}(R^{A},R^{B},0,0,0,0)\,\,.$ (1)
@@
$p_{\textit{reported}}^{AB}\mathrel{\mathop{:}}=\lim_{O^{A}\to 0}\frac{R^{B}}{R^{A}-O^{A}}=\frac{R^{B}}{R^{A}}\,\,.$ (5)
@@
Since trading fees are not added to the pool (as defined in Section 3.2), the product of $R^{A}$ and $R^{B}$ remains constant after each trade (Equation 2). Then, arbitrageurs keep the ratio of $R^{A}$ and $R^{B}$ equal to $\frac{p^{A}}{p^{B}}$ (Equation 6). Therefore, $R^{A}$ and $R^{B}$ remain the same after the trade and arbitrage.
4.3 CPMM Trading fee
AMMs charge a trading fee for each trade operation, which is paid by the trader. These trading fees form the revenue of liquidity providers. In CPMMs, the trading fee is a constant fraction ${1-\gamma\in[0,1]}$ of input tokens [5]. This is achieved by selecting the trading function $\Phi_{\textit{CPMM}}(R^{A},R^{B},I^{A},O^{A},I^{B},O^{B})$ as [5]
$\Phi_{\textit{CPMM}}(R^{A},R^{B},I^{A},O^{A},I^{B},O^{B})\mathrel{\mathop{:}}=\\ (R^{A}+\gamma I^{A}-O^{A})\times(R^{B}+\gamma I^{B}-O^{B})\,\,.$

### v2

@@ Abstract
Automated Market Makers (AMMs) are a cornerstone of decentralized finance (DeFi) blockchain-based platforms. They enable direct exchange of virtual tokens: Traders exchange tokens with the AMM, paying a fee; liquidity comes from liquidity providers, paid by those fees. Despite growing demand, the performance of AMMs is limited. State-of-the-art blockchain platforms allow for parallel execution of transactions. However, we show that AMMs do not enjoy these gains since their operations are not parallelizable.
We present SAMM, an AMM comprising multiple independent shards. All shards operate in the same chain, but they allow for parallel execution as each is independent. The challenge is that traders are incentivized to split each trade among all AMMs in existing designs, leading to lower throughput. SAMM addresses this issue with a novel design of the trading fees. Traders are incentivized to use only a single smallest shard. We show that all Subgame-Perfect Nash Equilibria (SPNE) fit the desired behavior: Liquidity providers balance the liquidity among all shards, so the system converges to the state where trades are evenly distributed, overcoming destabilization attacks.
Evaluation in the Sui and Solana blockchains shows that SAMM improves throughput by 5 and by 16, respectively, approaching their limit. SAMM is directly deployable, allowing trading at scale for individuals and DeFi applications.
I Introduction
Decentralized platforms like Ethereum [53] enable users to mint tokens and transfer them among each other. They can also transfer tokens to and receive tokens from stateful programs called smart contracts. Decentralized Finance (DeFi) encompasses a variety of financial smart contracts. Their users issue transactions (txs) to generate, loan, and exchange tokens. Automated Market Makers (AMMs) are a cornerstone of the DeFi ecosystem [23, 24]. They enable users to exchange token pairs by maintaining liquidity pools: tokens of both types supplied by other users serving as liquidity providers. The demand for AMMs grows rapidly: The prominent Uniswap [59, 4, 3] exchanged \$1 trillion in its first 42 months of operation and then an additional \$1 trillion within only 24 months [33]. If the current trend continues (Appendix A), by 2029 demand will surpass $200$ transactions per second (tps).
However, AMM throughput is limited due to the limits of the underlying blockchain. For example, the Uniswap AMM was first deployed in Ethereum, but its demand has surpassed Ethereum’s capacity (Appendix A). This leads users to defer some workload to expansion systems called layer-2 protocols [35, 2]; but this entails significant additional security assumptions. [^1] Blockchains are secured by financially-invested nodes. A layer-2 protocol relies on an additional set of nodes, with significantly less at stake [17]. It is thus vulnerable to cheaper attacks, at least in terms of fairness and progress. Moreover, insufficient AMM capacity leads to longer queues for AMM trades, exacerbating incentive-based vulnerabilities like sandwich attacks [60, 31].
@@ I Introduction
Figure 1: Sui trade latency as a function of demand with $n$ SAMM shards.
Figure 2: Solana trade latency as a function of demand with $n$ SAMM shards.
@@
In this work, we address the throughput limitation of AMMs by using multiple AMM instances called shards. We model the system (§III) as a set of AMM shards and rational users of two kinds. Traders purchase tokens, they use the available AMMs and pay fees as required, aiming to minimize their expenses. Liquidity providers deposit tokens into AMMs and earn fees based on their contribution.
@@
The shards are AMMs based on the standard Constant Product Market Maker (CPMM) contract (§IV): Roughly, the contract maintains the product of the two tokens constant after each trade. Thus, purchasing a larger amount of a token increases its unit cost, an effect called slippage.
We present SAMM (§V), an AMM protocol that uses multiple shards, all of which are AMM smart contracts operating on the same blockchain. But the decentralized nature of the setting raises a challenge. Ideally, all shards should be balanced, with the same liquidity; and traders should randomly select a shard to trade. However, traders and liquidity providers, acting rationally, can deviate from this desired behavior. Traders might split trades across multiple shards, multiplying their overhead. Liquidity providers could cause imbalances in shard sizes, making some shards less attractive to traders and reducing parallelism. Additionally, attackers might deliberately disrupt the system by destabilizing the shard balance. Existing AMM designs fail to address these issues. To overcome this, SAMM diverges from the traditional approach of using a fixed-ratio fee. Instead, it employs a trading fee function that incentivizes traders to use the smallest shard.
@@
The model gives rise to a game (§VI) played among the users. In each step, either a liquidity provider adds liquidity to a subset of the shards, or a trader executes a trade using a subset of the shards. We assume myopic liquidity provider behavior, reducing the analysis to a Stackelberg game [50] where the liquidity provider adds liquidity to maximize her revenue from a subsequent trade. We observe that naively using a set of independent CPMMs results in all trades being split among all CPMMs, increasing system overhead without improving the satisfied demand. Our analysis of SAMM shows that, indeed, in all best responses, traders use one of the smallest shards. This implies that when shards are balanced, traders will randomly select a shard to trade, allowing for maximal parallelization. In addition, we prove that shards converge (e.g., following intentional destabilization) to a balanced state in all subgame-perfect Nash Equilibria (SPNE) and then maintain balance.
@@
To evaluate SAMM (§VII), we implement the protocol and deploy it in Sui and Solana local test networks. In Sui, the throughput increase is limited by the serial elements of Sui’s transaction processing, following Amdahl’s Law. In Solana, the throughput increases linearly with the number of shards up to a limit due to the blockchain’s constraints. SAMM achieves over a fivefold throughput increase in Sui and a sixteenfold increase in Solana, compared to a standard single-contract AMM. Figures 1 and 2 show that with more shards in Sui and Solana, SAMM achieves higher throughput (X axis) with low average latency (Y axis). Error bars show additional experiments, $\times$ marking failure due to overload. Improving the blockchains’ parallelism would allow for even better performance.
Finally, we confirm the effects of our incentive design through a simulation using real trade data (§VIII). Traders exhibit the behavior predicted by our game-theoretic analysis of the theoretical model. Furthermore, the results demonstrate that SAMM significantly enhances liquidity providers’ revenue with only a slight increase in traders’ costs, due to its improved throughput that enables more trades.
@@
In summary, our contributions are: (1) Identification of the blockchain performance bottleneck due to “hot” contracts, (2) SAMM: A sharded AMM contract improving scalability without layer-2 security assumptions, (3) generalization of the trading fee function of AMMs and a novel trading fee function to incentivize SAMM’s desired behavior and ensure security, (4) game-theoretic analysis showing Subgame-Perfect Nash Equilibrium (SPNE), (5) evaluation in Sui and Solana, demonstrating a fivefold (in Sui) and sixteenfold (in Solana) increase in throughput (up to the blockchain’s limit), and (6) simulation with real trade data confirming the effects of SAMM’s incentive design.
@@
These results hint at an upcoming challenge (§IX) in smart contract platform design: minimizing the serial elements of transaction processing. But SAMM can already be employed to scale AMM performance both for direct usage and as part of DeFi smart contracts.
II Related Work
@@
The introduction of the constant product market maker (CPMM) model by Uniswap v1 [59] set a new standard for AMMs, employing a liquidity pool and an algorithm designed to keep the product of the token balances constant. This approach enabled asset exchanges without relying on traditional order books. Subsequent iterations, Uniswap v2 [4] and v3 [3], further developed the CPMM model by improving the price oracle mechanism (reference of the token exchange rate) and the returns for liquidity providers, respectively. As a result, the CPMM algorithm has become a benchmark, with many AMMs adopting similar trading mechanisms [38, 49, 15].
@@
Angeris et al.[5] expanded the understanding of AMMs by delving into constant function market makers (CFMMs), demonstrating their utility as decentralized price oracles and broadening the CPMM model’s application. Following this, research has increasingly focused on trading utility maximization [6], advanced arbitrage techniques [60, 27, 7, 51], improving liquidity providers’ returns [36, 24], ensuring transaction privacy [14], eliminating Miner Extractable Value (MEV) for fair trades [13, 54, 12], and examining the synergy between blockchain-based AMMs and prediction market mechanisms [44]. To the best of our knowledge, previous work did not address AMM throughput scaling.
@@
III Model
We abstract away the blockchain details for our analysis and model the system as a set of participants (§III-A) interacting directly with AMMs (§III-B), exchanging tokens. The system progresses in discrete steps (§III-C) and there exists an external market used by arbitrageurs (§III-D) to arbitrage in our system. The model uses a generic AMM, which we will later instantiate based on previous work and with SAMM.
III-A Participants
There are two types of participants, liquidity providers and traders. Each liquidity provider holds some token A and token B. They aim to increase their holdings. Traders are either AB or BA, based on their goals. Each BA trader occasionally wishes to get a certain amount of token A, and vice versa for AB traders. They have sufficiently many tokens of the opposite type to complete their trade, but they aim to minimize the cost of obtaining the desired tokens. Both liquidity providers and traders can send and receive tokens to and from the smart contracts.
III-B Automated Market Makers
@@
  • Trade: The trader sends $I^{A}$ token A (resp., $I^{B}$ token B) to the contract and gets $O^{B}$ token B (resp., $O^{A}$ token A) from the contract.
@@
Following prior work [36, 24] we assume that the trading fee is directly paid to the liquidity providers. Although some practical AMMs (e.g., Uniswap v2[4]) reinvest the trading fees into the liquidity pool, allowing liquidity providers to withdraw more tokens than they deposited as utilities, the trading fee’s impact is negligible relative to the deposited amount: The average ratio of the output amount to the deposited amount is nominal (less than $0.036\%$ as we find in Appendix B), and automated market makers (AMMs) typically charge a low trading fee relative to the gross amount (e.g., $0.3\%$ in Uniswap), so the trading fee’s impact is negligible relative to the amount of deposited tokens.
III-C System State and Progress
There are $n$ independent AMM liquidity pools $\textit{pool}_{1},\textit{pool}_{2},\cdots\textit{pool}_{n}$ in the system. The system progresses in discrete steps $k=0,1,2,\cdots$ and is orchestrated by a scheduler. In each step, the scheduler randomly selects a participant and this participant executes transactions. It chooses a liquidity provider with probability $P_{lp}\geq 0$ and a trader with probability $P_{t}\geq 0$, where $P_{lp}+P_{t}=1$.
@@ III-C System State and Progress
The trader is either a BA trader aiming to obtain token A or an AB trader aiming to obtain token B. The probability of drawing an AB trader (resp., BA trader) is $P_{t}^{\textit{AB}}\geq 0$ (resp., $P_{t}^{\textit{BA}}\geq 0$), with $P_{t}^{\textit{AB}}+P_{t}^{\textit{BA}}=P_{t}$. To avoid repetition, we only show the case of BA traders, the expressions for AB traders are symmetric. The scheduler assigns the BA trader an amount of $b^{\textit{BA}}$ token A to obtain, following a random distribution $D^{\textit{BA}}$.
III-D External Market and Arbitrageurs
Following prior work [24, 36], we assume there is an external market providing the price of token A and token B, $p^{A}$ and $p^{B}$, respectively. These prices do not change due to trades; they serve as objective prices for token A and token B.
If the AMM sells tokens at a price lower than the external market, a principal can buy tokens from the AMM and sell them in the external market to make profits, or vice versa if the price is higher in the AMM; this is called an arbitrage. Previous work [36, 24, 12, 13] assumes active and rational arbitrageurs who can use the external market and arbitrage to maximize their utility. We follow the assumption [36] that there are immediate arbitrages without trading fees when the token price in the AMM is different from the external market.
III-E Our Goal
@@
IV Preliminaries: CPMM
The basis of SAMM is the Constant Product Market Maker AMM (CPMM, e.g. [3, 38, 49, 15]). We review its operation here. It uses a share-based solution to manage liquidity addition and removal operations (§IV-A) and keeps the product of the deposited amount of two tokens constant in trade operations (§IV-B). Liquidity providers earn revenue from the trading fees (§IV-C).
IV-A Liquidity Addition and Removal
Most CPMMs (e.g., [59, 4, 38, 41]) use a fungible share token to manage liquidity addition and removal. These tokens represent a liquidity provider’s share in the pool.
When liquidity providers add tokens to an AMM, they receive share tokens which signify their portion of the pool. To recall, $R^{A}$ and $R^{B}$ denote the amounts of token A and token B already deposited in the pool. Similarly, $I^{A}$ and $I^{B}$ represent the quantities of token A and token B that the liquidity provider contributes through a liquidity addition operation. Let $R^{S}$ represent the total amount of all share tokens distributed before the operation. The amount of share tokens acquired by the liquidity provider in this operation, $O^{S}$, is given by
@@ IV-A Liquidity Addition and Removal
The term $\min\left\{\frac{I^{A}}{R^{A}},\frac{I^{B}}{R^{B}}\right\}$ signifies the ratio of the input token to the deposited token. The $\min$ function serves to ensure that the ownership accurately reflects the liquidity provider’s contribution relative to the scarcer asset. It prevents situations where a liquidity provider inputs a large amount of a certain token to unfairly obtain a larger share of tokens in the pool.
Liquidity providers have the option to withdraw tokens from the pool with the liquidity removal operation, which takes the number of share tokens as input and outputs token A and token B. Let $I^{S}$ represent the amount of input share tokens, and $R^{S}$ denote the total amount of all share tokens referred to the pool before the execution. The amounts of token A and token B withdrawn are $O^{A}$ and $O^{B}$:
@@ IV-B CPMM Trades
$\Phi^{\textit{net}}_{\textit{CPMM}}(R^{A},R^{B},I^{A},O^{A},I^{B},O^{B})\coloneqq\\ (R^{A}+I^{A}-O^{A})\times(R^{B}+I^{B}-O^{B})\,\,.$
@@
$\Phi^{\textit{net}}_{\textit{CPMM}}(R^{A},R^{B},I^{A},O^{A},I^{B},O^{B})=\\ \Phi^{\textit{net}}_{\textit{CPMM}}(R^{A},R^{B},0,0,0,0)\,\,.$ (1)
@@
$p_{\textit{reported}}^{AB}\coloneqq\lim_{O^{A}\to 0}\frac{R^{B}}{R^{A}-O^{A}}=\frac{R^{B}}{R^{A}}\,\,.$ (5)
@@
Since trading fees are not added to the pool (as defined in Section III-B), the product of $R^{A}$ and $R^{B}$ remains constant after each trade (Equation 2). Then, arbitrageurs keep the ratio of $R^{A}$ and $R^{B}$ equal to $\frac{p^{A}}{p^{B}}$ (Equation 6). Therefore, $R^{A}$ and $R^{B}$ remain the same after the trade and arbitrage.
IV-C CPMM Trading fee
AMMs charge a trading fee for each trade operation, which the trader pays. These trading fees form the revenue of liquidity providers. In CPMMs, the trading fee is a constant fraction ${1-\gamma\in[0,1]}$ of input tokens [5]. This is achieved by selecting the trading function $\Phi_{\textit{CPMM}}(R^{A},R^{B},I^{A},O^{A},I^{B},O^{B})$ as [5]
$\Phi_{\textit{CPMM}}(R^{A},R^{B},I^{A},O^{A},I^{B},O^{B})\coloneqq\\ (R^{A}+\gamma I^{A}-O^{A})\times(R^{B}+\gamma I^{B}-O^{B})\,\,.$